# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ VI Benefits and Applications of BD-RIS
Thanks to the flexible interconnections between elements to support various architectures and modes, BD-RIS has multiple benefits, such as enhancing channel gain, increasing transmission quality, enlarging coverage, etc. The appealing benefits of BD-RIS have also spawned an explosion of application-oriented works together with other promising 5G/6G techniques, such as advanced multiple access techniques, wireless sensing, and wireless power transfer (WPT). In this section, we summarize the key benefits and emerging applications of BD-RIS.
VI-A Benefits
VI-A.1 Boosting Received Power and Rates
Compared to (lossless) D-RIS which can only manipulate the phase shift of the diagonal entries in the scattering matrix, BD-RIS has higher flexibility in manipulating the amplitude and phase shift of both diagonal and off-diagonal entries in its scattering matrix. This flexibility boosts system performance in various scenarios.
Specifically, for the SISO system where the wireless channels are modelled as i.i.d. Rayleigh fading, [26] has derived the scaling law of the received signal power as a function of the number of BD-RIS elements $M$. The power gain of the group/fully-connected BD-RIS over D-RIS is quantified by
$G^{\mathsf{group-conn}}=\frac{\bar{M}+\frac{M-\bar{M}}{\bar{M}^{2}}\left(\frac{\Gamma(\bar{M}+\frac{1}{2})}{\Gamma(\bar{M})}\right)^{4}}{1+(\bar{M}-1)\Gamma^{4}(1.5)},$ (108)
where $\Gamma(\cdot)$ denotes the gamma function. The limits of the power gain when $M\rightarrow\infty$ is thus given by
$\lim_{M\rightarrow\infty}G^{\mathsf{group-conn}}=\frac{1}{\bar{M}^{2}}\left(\frac{\Gamma(\bar{M}+\frac{1}{2})}{\Gamma(\bar{M})\Gamma(1.5)}\right)^{4}.$ (109)
For fully-connected BD-RIS with $\bar{M}=M$, the limit of the power gain can be as high as $\frac{1}{\Gamma^{4}(1.5)}=\frac{16}{\pi^{2}}\approx 1.62$, which has also been illustrated in Section II. This indicates that BD-RIS can increase the received power by up to 62% in SISO systems compared to D-RIS.
For more general multi-user MISO systems, [54, 56, 57] have shown significant sum-rate gains of fully/group-connected BD-RIS over D-RIS. A representative example is provided in Fig. 25, where the fully-connected BD-RIS can increase the sum-rate by around 43% (under specific parameter settings) compared to D-RIS.
Figure 25: Sum-rate versus transmit power for different BD-RIS architectures. Channels through BD-RIS follow Rician fading with a Rician factor 2 dB. The transmitter-RIS distance is set as 50 m and the RIS-user distance is set as 2.5 m. ($N=K=4$, $M=32$, $\bar{M}=4$ for the group-connected architecture) [57].
VI-A.2 Enabling Architectures with Low Complexity to Achieve High Performance
The flexibility provided by the interconnections between BD-RIS elements enables various circuit topology designs to support different architectures as illustrated in Sections III-C and III-D. More importantly, this flexibility enables architectures with least hardware complexity (i.e., number of tunable components) to achieve optimal performance in various systems.
For the MISO system, [40] has theoretically proven that the least circuit complexity for MISO optimal BD-RIS is
$C_{\mathsf{MISO}}^{\mathsf{opt}}=2M-1.$ (110)
This condition essentially forms a tree-connected architecture, and two representative examples are arrowhead BD-RIS and tridiagonal BD-RIS. Based on this conclusion, [111] has analyzed the Pareto frontier between performance and circuit complexity for a SISO system. Results in Fig. 26 show that tree-connected architecture can reach the performance achieved by fully-connected BD-RIS with the lowest possible complexity, thus achieving the best performance-complexity trade-off.
Figure 26: Pareto frontier between performance and circuit complexity including single-, group-, fully-, forest-, and tree-connected BD-RIS, with different group sizes $\bar{M}$ labeled on two sides of the curve. Channels through BD-RIS follow i.i.d. Rayleigh fading.
For more general multi-user MIMO systems, [41] has theoretically proved that the least circuit complexity for multi-user MIMO optimal BD-RIS is
$C_{\mathsf{MU-MIMO}}^{\mathsf{opt}}=\min\{D,\frac{M}{2}\}\left(2M-2\min\{D,\frac{M}{2}\}+1\right),$ (111)
where $D=\min\{\sum_{k\in\mathcal{K}}N_{k},N\}$ is the DoF of the multi-user MIMO channel. Two representative examples which can reach such condition are band-connected and stem-connected BD-RIS with band and stem width $q=2\min\{D,\frac{M}{2}\}-1$. Simulation results in Fig. 27 show the Pareto frontier between performance and circuit complexity achieved by single-, group-, fully-, stem-, and band-connected BD-RIS-aided multi-user MISO systems, indicating that band/stem-connected BD-RIS with proper band/stem width $q$ can reach the performance achieved by fully-connected BD-RIS, but with much less circuit complexity.
Figure 27: Pareto frontier between performance and circuit complexity including single-, group-, fully-, stem-, and band-connected BD-RIS, with different group sizes $\bar{M}$ and stem/band width $q$ labeled on two sides of the curves. Channels through BD-RIS follow Rician fading with a Rician factor 2 dB. The transmitter-RIS distance is set as 50 m and the RIS-user distance is set as 2.5 m. ($N=K=4$, $M=64$, $P=10$ dBm) [41].
VI-A.3 Enabling Flexible Modes to Achieve Highly-Directional Wireless Coverage
The inter-element connections in BD-RIS not only boost the channel strength and system performance, but also provide possibility to flexibly arrange the locations and orientations of elements. This leads to different modes with enlarged wireless coverage.
The hybrid transmitting and reflecting mode has been proposed in [36] by back to back placing every two elements and interconnect them to construct a cell. This arrangement enables signals impinging on one sector of BD-RIS to be partially reflected from the same sector and partially transmitted to the other sector, achieving full-space coverage. Results in Fig. 28 demonstrate that hybrid BD-RIS always outperforms transmitting/reflecting BD-RIS since the former could simultaneously serve users from both sectors while half of the users are blocked when using transmitting/reflecting BD-RIS.
Figure 28: Sum-rate for a BD-RIS-aided multi-user MISO system versus the number of BD-RIS cells. “CW-FC”, “CW-GC”, and “CW-SC”, respectively, refer to cell-wise fully-, group-, and single-connected. Channels through BD-RIS follow Rician fading with a Rician factor 5 dB. Half users are located within the coverage of the reflecting sector and the other half are located within the coverage of the transmitting sector. The transmitter-RIS distance is set as 50 m and the RIS-user distance is set as 2.5 m. ($N=K=6$, $G=8$ for cell-wise group-connected architecture, $P=5$ dBm) [36].
Figure 29: Sum-rate for a BD-RIS-aided multi-user MISO system versus the number of BD-RIS elements. Channels through BD-RIS follow Rician fading with a Rician factor 0 dB. Among all $K$ users, every $K/L$ users are located within the coverage of one sector. The transmitter-RIS distance is set as 100 m and the RIS-user distance is set as 10 m ($N=K=6$, $G=6$ for cell-wise group-connected architecture, $P=30$ dBm) [38].
In addition to the hybrid mode proposed in [36], a more flexible multi-sector mode has been proposed in [38], where every $L$, $L\geq 2$ elements are located at the edge of a polygon and interconnected with each other to construct a cell. As such, signals impinging on one sector can be partially reflected within the same sector and partially scattered toward other sectors to enable full-space coverage. More importantly, this element arrangement enables the use of antennas with narrower beamwidth and thus higher gains, thereby providing additional enhanced channel gains. Results in Fig. 29 demonstrate that multi-sector BD-RIS with a larger number of sectors achieves better performance, thanks to the higher antenna gains provided by more directional antennas.
VI-A.4 Providing Orders of Magnitude Gains over D-RIS in Distributed Deployments
In most existing literature, an RIS has been regarded as an antenna array whose inter-element spacing is comparable with (typically smaller than) the wavelength. This setting results in a localized RIS whose elements are localized in a specific site. The localized RIS has been widely studied in far-field scenarios, where the wireless channels related to each RIS element share approximately the same large-scale fading, as shown in Fig. 30(a). In this sense, the gain of BD-RIS over D-RIS mainly comes from the better exploration of small-scale fading effects. To understand if the joint effects of small- and large-scale fading in wireless channels can provide additional flexibility to be captured by BD-RIS architectures, [91] has proposed the concept of distributed RIS. In the distributed RIS, elements are distributed over a wide region with inter-element spacing much larger than the wavelength, such that the large-scale fading between transmitter/receiver and each RIS element can be quite different, as shown in Fig. 30(b).
To quantify the benefit of distributed RIS, [91] has derived the gain of distributed over localized fully-connected BD-RIS in SISO systems under Rayleigh fading channels, which is bounded by
where $d_{R}$ and $d_{T}$, respectively, denote the distance between the localized RIS and the receiver/transmitter; $\mathbf{d}_{R}\in\mathbb{R}^{M\times 1}$ and $\mathbf{d}_{T}\in\mathbb{R}^{M\times 1}$, respectively, denote the distance between all $M$ elements of the distributed RIS and the receiver/transmitter; $a$ denotes the path-loss exponent. The bounds in (112) indicate that to guarantee $G^{\mathsf{dis}}>1$, a sufficient condition is to have $d_{R}d_{T}>\sqrt[a]{M^{2}}\min\{\mathbf{d}_{R}\}\min\{\mathbf{d}_{T}\}$ and a necessary condition is to have $d_{R}d_{T}>\min\{\mathbf{d}_{R}\}\min\{\mathbf{d}_{T}\}$, both of which are mild and can be easily achieved. This is also numerically evaluated in Fig. 31(a), showing that $G^{\mathsf{dis}}$ grows exponentially with $a$ and reaches as high as several orders of magnitude. In addition, the impact of locations of the receiver is reported in Fig. 31(b), showing that the distributed arrangement has the most significant performance benefit when the receiver is far from the localized BD-RIS while being close to the distributed BD-RIS.
(a) Gain of distributed over localized BD-RIS $G^{\mathsf{dis}}$
(b) Gain of distributed over localized BD-RIS $G^{\mathsf{dis}}$
Figure 31: $G^{\mathsf{dis}}$ (in dB) for different values of path-loss exponent, number of RIS elements, and locations of receiver [91].
VI-A.5 Enabling Simultaneously Optimal Transmissions for Uplink and Downlink with Non-Reciprocal Architectures
One key property of non-reciprocal BD-RIS is that its admittance/impedance/scattering matrices are asymmetric, which naturally break the uplink-downlink reciprocity of the wireless channels. For example, given an uplink BD-RIS-aided MISO channel $\mathbf{h}_{\mathsf{up}}=\mathbf{H}_{TI}\mathbf{\Theta}\mathbf{h}_{IR}$ as shown in (81), the downlink channel $\mathbf{h}_{\mathsf{down}}=\mathbf{h}_{RI}\mathbf{\Theta}\mathbf{H}_{IT}=\mathbf{h}_{IR}^{\mathsf{T}}\mathbf{\Theta}\mathbf{H}_{TI}^{\mathsf{T}}$ is not equal to the transpose of the uplink channel, i.e., $\mathbf{h}_{\mathsf{down}}\neq\mathbf{h}_{\mathsf{up}}^{\mathsf{T}}$ since $\mathbf{\Theta}\neq\mathbf{\Theta}^{\mathsf{T}}$. This property provides unique benefits to scenarios where uplink and downlink transmissions behave differently, such as full-duplex systems [112].
To visualize the benefit of applying non-reciprocal BD-RIS in full-duplex systems, [113] has derived the general RIS-aided full-duplex system model, and the theoretical conditions for non-reciprocal BD-RIS to simultaneously maximize the received powers of the signal of interest in the uplink and downlink. Results in the top two figures of Fig. 32 show that, when the uplink and downlink users are not aligned, both the impinging and reflected beams of non-reciprocal BD-RIS can exactly point to the directions of the signal of interest with the maximum power, while those of reciprocal BD-RIS fail to point to the downlink user. Meanwhile, the bottom two figures in Fig. 32 show that, only when the uplink and downlink users are aligned, both the impinging and reflected beams of reciprocal and non-reciprocal BD-RIS can exactly point to the directions of signal of interest. Due to the unique property of non-reciprocal BD-RIS, if a wave hits the non-reciprocal BD-RIS from one direction, the surface behaves differently than if it hits from the opposite direction. This finally enables an uplink user and a downlink user at different locations to optimally communicate with the same full-duplex base station. While [113] primarily focuses on a single-antenna scenario, the benefit of non-reciprocal BD-RIS in full-duplex systems has been recently shown in more general multi-user multi-antenna scenarios [114].
Figure 32: The impinging and reflected beam patterns of reciprocal and non-reciprocal BD-RISs. The base station (marked as “BS”) is located at $\frac{\pi}{6}$; the downlink user is located at $\frac{\pi}{2}$; and the uplink user is located at $\frac{2\pi}{3}$ in the top figures and at $\frac{\pi}{2}$ in the bottom figures [113].
VI-A.6 Providing Enhanced Gains over D-RIS in Dual-Polarized Systems
Existing BD-RIS literature has focused on uni-polarized systems for the ease of analysis. However, modern MIMO systems use dual-polarized antenna arrays to have more antennas within limited space to provide more beamforming gains [115]. Dual-polarized D-RIS-aided systems have been studied in [116, 117], while it remains unknown if BD-RIS still has performance gains over D-RIS in dual-polarized systems. Taking into account this practical consideration, [118] has modeled and analyzed a dual-polorized BD-RIS-aided SISO system, where the BD-RIS has half vertically polarized elements and half horizontally polarized elements, and the receiving and transmitting antennas can have the same or opposite polarizations, as illustrated in Fig. 33.
When the transmitter and receiver have the same polarizations, as illustrated in Fig. 33(a), the limit of the power gain of dual-polarized fully-connected BD-RIS over D-RIS remains the same as the uni-polarized case [26]. That is, dual-polarized BD-RIS can increase the received power in SISO by up to 62% compared to dual-polarized D-RIS. When the transmitter and receiver have opposite polarizations, the limit of the power gain of dual-polarized fully-connected BD-RIS over D-RIS varies with channel fadings, and is expressed as [118]
where $10<\chi<$ 1 denotes the inverse of the cross-polar discrimination. This analytical result indicates that BD-RIS can offer a gain $G^{\mathsf{dual-polarized}}>1$ for both Rayleigh fading and LoS channels. To visualize the impact of $\chi$ on performance gain, Fig. 34 illustrates the gain of BD-RIS over D-RIS as a function of $\chi$. These analytical results show that BD-RIS has more significant benefits over D-RIS with a smaller $\chi$, as small values of $\chi$ increase disparity between the channel entries.
Figure 34: Gain of BD-RIS over D-RIS $\lim_{M\rightarrow\infty}G^{\mathsf{dual-polarized}}$ as a function of $\chi$ when the transmitter and receiver have the same/opposite polarization and with Rayleigh/LoS channels [118].
VI-B Emerging Applications
The appealing benefits of BD-RIS have triggered many emerging applications, as discussed below and summarized in Table VIII.
VI-B.1 BD-RIS for Channel Shaping
@@
Most existing BD-RIS literature focuses on specific metrics, such as channel gain, sum-rate, transmit power consumption, to evaluate the performance benefit of BD-RIS, while a fundamental question which has not been answered is, to what extent a BD-RIS can shape the wireless channels (in terms of their singular values). To answer this question, [50] has theoretically proven that BD-RIS is able to reach a larger dynamic range of channel singular values than D-RIS, followed by numerical verifications supported by algorithm designs. Results in [50] have shown that the capability of more flexible channel shaping in BD-RIS can help to increase power gain and achievable rates. Further, [119, 120] have recently shown that BD-RIS can achieve better interference nulling/minimization to MIMO interference channels, which can be explained by the potential of BD-RIS to reduce the DoF.
VI-B.2 BD-RIS Aiding Channel Reciprocity Attack and Physical Layer Security
@@
According to microwave engineering, conventional wireless channels are naturally reciprocal, in the sense that the uplink channel is the transpose of the downlink channel. However, as explained in the previous subsection, introducing non-reciprocal BD-RIS in the wireless propagation environment can break the uplink-downlink reciprocity of wireless channels. From the perspective of physical layer security, this property of non-reciprocal BD-RIS can be beneficial to support simultaneously optimal uplink and downlink transmission when downlink and uplink legitimate users are not aligned, as illustrated in Fig. 32. It can also be helpful to enhance the secrecy rate and secrecy outage probability [121] in the presence of eavesdroppers. From an attacker’s perspective, this property can still be useful since a non-reciprocal BD-RIS can be potentially deployed to maliciously degrade the downlink network performance [122, 123].
@@
[50] Channel Shaping Non-Reciprocal Reflecting Channel Gain & Sum-Rate Max.
[119] Fully/Group-Connected Sum-Rate Max. & Interference Nulling
[120] Interference Min.
[121] Physical Layer Security Non-Reciprocal Secrecy Rate Max. & Secrecy Outage Prob. Min.
[122, 123] Reciprocity Attack Sum-Rate Min.
[124] THz Commun. Hybrid Sum-Rate Max.
[125] Massive MIMO Reflecting Max. Min. Rate
[126] Fully/Group-Connected Hybrid Bit Error Rate Min. & Rate. Max.
[127] Non-Reciprocal Sum-Rate Max.
[128] SIM Fully/Group-Connected Channel Gain Max.
[129] NOMA Fully/Group-Connected Reflecting Sum-Rate Max.
[130] Non-Reciprocal Max. Min. Rate
[131, 132, 133] RSMA Fully/Group-Connected Sum-Rate Max.
[134] Max. Min. Rate
[135] Group-Connected Multi-Sector Ergodic Sum-Rate Max. (Imperfect CSI)
[136] UAV Non-Reciprocal Reflecting Sum-Rate Max.
[137] Min. Max. Computational Time
[138] Fully/Group-Connected Zero Secrecy Rate Prob. Min.
[139] NTN Non-Reciprocal Secrecy Rate Max.
[140] Rate Max.
[141] Wireless Sensing Group-Connected Multi-Sector Mean Square Error Min.
@@
[143] ISAC Non-Reciprocal Hybrid Max. Min. Signal-to-Clutter-Plus-Noise Ratio
[144] Fully-Connected Reflecting Weighted Sum of Radar and Commmun. SNRs Max.
[145] Power Min.
[146] Sum-Rate (Throughput) Max.
@@
[148] Outage Probability Analysis
[149, 150] WPT Direct Current Harvested Power Max.
[151] SWIPT Sum-Rate Max. & Harvested Power Max.
[152] MEC Non-Reciprocal Hybrid No. of Completed Task Bits Max.
VI-B.3 BD-RIS in Terahertz Communications and Massive MIMO
@@
The need for more spectrum resources and higher frequency bands, such as Terahertz frequency bands, to provide higher spectrum efficiency and support emerging applications is one important trend for future networks. However, one bottleneck of Terahertz communications is the severe path loss in wireless channels, which makes the transmission vulnerable to blockages [153]. BD-RIS has been shown to be a reliable solution to Terahertz communications, which can be useful to bypass blockages, increase coverage and system rates [124].
@@
To compensate for the severe path loss in higher frequency bands, serve a large number of users in the same time-frequency resource, and support increasing connectivity in 6G, another promising technique is massive MIMO [154]. However, the high spectrum efficiency of massive MIMO is achieved at the cost of numerous RF chains, leading to energy inefficiency in the network. To reduce the required number of RF chains while maintaining satisfactory spectrum efficiency, [125, 126] have recently proposed to integrate BD-RIS within the radome of a transmitter as an auxiliary passive array, enabling a small-dimensional active antenna array at the transmitter to provide highly-directional beams reconfigured by BD-RIS. While [125, 126] focused on single-transmitter scenarios, [127] has considered a more general cell-free network and verified the performance benefits provided by BD-RIS.
VI-B.4 BD-RIS and Stacked Intelligent Metasurface (SIM)
Stacked intelligent surface (SIM) [155] consists in stacking multiple layers of RIS or metasurfaces to provide more flexible wave manipulations. The SIM, usually with a large dimension to provide sufficient wave manipulation capability, is typically integrated with transmitters/receivers having only a small number of active antennas. In this way, the requirement of cost- and energy-hungry devices, such as RF chains, analog-to-digital converters and digital-to-analog converters, is much reduced. While SIM layers have been implemented using D-RIS [155], [128] has recently shown that SIM layers can be implemented using BD-RIS. Specifically, the general modeling of SIM aided communication systems has been first established, followed by physically consistent assumptions and simplifications. Theoretical and numerical results have demonstrated that 1-layer SIM implemented with BD-RIS (tree-connected) is sufficient to achieve performance upper-bound in SISO systems, while multi-layer SIM implemented with D-RIS is suboptimal.
VI-B.5 BD-RIS with Advanced Multiple Access
Multiple access refers to techniques that make use of resource dimensions, such as time, frequency, space, and power, to serve multiple users. Among various enabling multiple access techniques, non-orthogonal multiple access (NOMA) has been actively studied in recent years due to its benefits in exploiting the available resources more efficiently with the aid of superposition coding at the transmitter and successive interference cancellation (SIC) at the receiver [156]. To further boost system performance, the interaction between NOMA and BD-RIS has been studied in [129, 130], showing that better user fairness can be achieved.
In addition, rate-splitting multiple access (RSMA) is a novel and general framework for non-orthogonal transmissions, softly including NOMA as an extreme case [157] and moving beyond that to provide enhanced spectral efficiency, energy efficiency, coverage, fairness, reliability, etc. Given that channel acquisition of RIS-aided wireless systems is generally challenging while RSMA has the benefit of being robust to CSI imperfections, the integration of the two has emerged [131, 132, 134, 133, 135]. Specifically, [131, 132, 134, 133] have focused mainly on perfect CSI and have shown the benefits of integrating BD-RIS and RSMA in increasing spectral efficiency, fairness rate, and fairness energy efficiency. [135] has further shown that the integration of the two can still increase system sum-rate when only imperfect CSI is available at the transmitter.
VI-B.6 BD-RIS Aiding Unmanned Aerial Vehicle (UAV) and Non-Terrestrial Networks (NTN)
Unmanned Aerial Vehicle (UAV) has gained attention for 5G and 6G networks due to its advantages in extending coverage, easing the deployment, and supporting controllable mobility [158]. These advantages of UAVs can be further strengthened by cooperating with BD-RIS. Specifically, [136] studied the transmission strategy of a ground-based BD-RIS in a network consisting of multiple UAVs, showing that the sum-rate performance has been much improved compared to scenarios adopting D-RIS and conventional scenarios without RIS. Focusing on the same system setting but in the presence of flying eavesdroppers, [138] proposed multi-user scheduling schemes to improve the physical layer security. In addition, [137] proposed to colocate BD-RIS and UAVs to minimize the computation latency and UAV hovering time.
In addition to UAV-enabled systems, BD-RIS can also provide performance benefits in more general non-terrestrial networks (NTN) typically realized by satellites, high-altitude platform stations (HAPS), and UAVs [159]. Results in [139, 140] have shown that having a BD-RIS mounted UAV [139] or a BD-RIS mounted HAPS [140] as a secondary transmitter in a cognitive radio enabled NTN can help increase the secrecy rate and spectral efficiency.
VI-B.7 BD-RIS with Sensing and ISAC

### v2

@@ VII Applications of BD-RIS
The appealing benefits of BD-RIS have spawned an explosion of application-oriented works together with other promising 5G/6G techniques, such as advanced multiple access techniques, sensing, and WPT. In this section, we summarize the emerging applications of BD-RIS, as detailed below and summarized in Table VIII.
VII-A BD-RIS for Communications
VII-A.1 BD-RIS for Channel Shaping
@@
Most existing BD-RIS literature focuses on specific metrics, such as channel gain, sum-rate, transmit power consumption, to evaluate the performance benefit of BD-RIS, while a fundamental question which has not been answered is, to what extent a BD-RIS can shape the wireless channels (in terms of their singular values). To answer this question, [zhao2024channel] has theoretically proven that BD-RIS is able to reach a larger dynamic range of channel singular values than D-RIS. Results in [zhao2024channel] have shown that this capability of BD-RIS can help to increase power gain and achievable rates. Further, [yahya2024beyond, santamaria2025interference] have recently shown that BD-RIS can achieve better interference nulling/minimization to MIMO interference channels, which can be explained by the potential of BD-RIS to reduce the DoF.
VII-A.2 BD-RIS Aiding Channel Reciprocity Attack and Physical Layer Security
@@
According to microwave engineering, conventional wireless channels are naturally reciprocal, in the sense that the uplink channel is the transpose of the downlink channel. However, as explained in the previous subsection, introducing non-reciprocal BD-RIS in the wireless propagation environment can break the uplink-downlink reciprocity of wireless channels. From the perspective of physical layer security, this property of non-reciprocal BD-RIS can be beneficial to support simultaneously optimal uplink and downlink transmission when downlink and uplink legitimate users are not aligned, as illustrated in Fig. 22. Non-reciprocal BD-RIS can also be helpful to enhance the secrecy rate and secrecy outage probability [agarwal2023enhanced] and enable a wireless circulator for one-way secure communications [liu2025secure] in the presence of eavesdroppers. From an attacker’s perspective, this property can still be useful since a non-reciprocal BD-RIS can be potentially deployed to maliciously degrade the downlink network performance [wang2024channel, wang2024beyond].
@@
[zhao2024channel] Channel Shaping Non-Reciprocal Reflecting Channel Gain & Sum-Rate Max.
[yahya2024beyond] Fully/Group-Connected Sum-Rate Max. & Interference Nulling
[santamaria2025interference] Interference Min.
[agarwal2023enhanced] Physical Layer Security Non-Reciprocal Secrecy Rate Max. & Secrecy Outage Prob. Min.
[liu2025secure] Sum-rate Max.
[wang2024channel, wang2024beyond] Reciprocity Attack Sum-Rate Min.
[mahmood2024enhancing] THz Commun. Hybrid Sum-Rate Max.
[mishra2023transmitter] Massive MIMO Reflecting Max. Min. Rate
[raeisi2025modern] Fully/Group-Connected Hybrid Bit Error Rate Min. & Rate. Max.
[li2025beamforming] Non-Reciprocal Sum-Rate Max.
[zhang2024full] Wireless Sensing Group-Connected Multi-Sector Mean Square Error Min.
@@
[wang2024dual] ISAC Non-Reciprocal Hybrid Max. Min. Signal-to-Clutter-Plus-Noise Ratio
[esmaeilbeig2024beyond] Fully-Connected Reflecting Weighted Sum of Radar and Commmun. SNRs Max.
[guang2024power] Power Min.
[liu2024enhancing] Sum-Rate (Throughput) Max.
@@
[nguyen2025beyond] Outage Probability Analysis
[azarbahram2025beyond, azarbahram2025beamforming] WPT Direct Current Harvested Power Max.
[hua2024cell] SWIPT Sum-Rate Max. & Harvested Power Max.
[nerini2024physically] SIM Fully/Group-Connected Hybrid Channel Gain Max.
[zhang2024beyond] NOMA Fully/Group-Connected Reflecting Sum-Rate Max.
[agarwal2025fairness] Non-Reciprocal Max. Min. Rate
[fang2022fully, kim2023group, khisa2024gradient] RSMA Fully/Group-Connected Sum-Rate Max.
[soleymani2023optimization] Max. Min. Rate
[li2024synergizing] Group-Connected Multi-Sector Ergodic Sum-Rate Max. (Imperfect CSI)
[huroon2023optimized] UAV Non-Reciprocal Reflecting Sum-Rate Max.
[mahmood2023joint] Min. Max. Computational Time
[lin2024securing] Fully/Group-Connected Zero Secrecy Rate Prob. Min.
[khan2025enhancing] NTN Non-Reciprocal Secrecy Rate Max.
[khan2025beyondnew] Rate Max.
[qin2025joint] MEC Non-Reciprocal Hybrid No. of Completed Task Bits Max.
VII-A.3 BD-RIS in Terahertz Communications and Massive MIMO
@@
The need for more spectrum resources and higher frequency bands, such as Terahertz frequency bands, to provide higher spectrum efficiency and support emerging applications is one important trend for future networks. However, one bottleneck of Terahertz communications is the severe path loss in wireless channels, which makes the transmission vulnerable to blockages [akyildiz2022terahertz]. BD-RIS has been shown to be a reliable solution to Terahertz communications, which can be useful to bypass blockages, increase coverage and system rates [mahmood2024enhancing].
@@
To compensate for the severe path loss in higher frequency bands, serve numerous users in the same time-frequency resource, and support increasing connectivity in 6G, another promising technique is massive MIMO [bjornson2017massive]. However, the high spectrum efficiency of massive MIMO is achieved at the cost of numerous RF chains, leading to energy inefficiency in the network. To reduce the required number of RF chains while maintaining satisfactory spectrum efficiency, [mishra2023transmitter, raeisi2025modern] have recently proposed to integrate BD-RIS within the radome of a transmitter as an auxiliary passive array, enabling a small-dimensional active antenna array at the transmitter to provide highly-directional beams reconfigured by BD-RIS. While [mishra2023transmitter, raeisi2025modern] focused on single-transmitter scenarios, [li2025beamforming] has considered a more general cell-free network and verified the performance benefits provided by BD-RIS.
VII-B BD-RIS for Sensing and ISAC